# Antonio Ramírez-de-Arellano Dept. of Computer Science and Artificial Intelligence

### v1

@@ Abstract
In the present work, we further study the computational power of virus machines (VMs in short). VMs provide a computing paradigm inspired by the transmission and replication networks of viruses. VMs consist of process units (called hosts) structured by a directed graph whose arcs are called channels and an instruction graph that controls the transmissions of virus objects among hosts. The present work complements our understanding of the computing power of VMs by introducing normal forms; these expressions restrict the features in a given computing model. Some of the features that we restrict in our normal forms include (a) the number of hosts, (b) the number of instructions, and (c) the number of virus objects in each host. After we recall some known results on the computing power of VMs we give our normal forms, such as the size of the loops in the network, proving new characterisations of family of sets, such as the finite sets, semilinear sets, or NRE.
@@ 1 Introduction
This paper contributes the following to the study of virus machines and their computing power. Normal forms, some of which are optimal bounds, for VMs are provided in the following sense: (a) providing characterisations (previously were inclusions) for generating families of finite sets; (b) showing new characterisations for finite sets of numbers using restrictions on the number of required hosts, instructions, or viruses; and (c) new characterisations are also given for singleton sets of numbers and some linear progressions or combinations. We also consider new restrictions: limiting or not the host or instruction graphs to be a tree graph, that is, an acylcic directed graph. Moreover, the instruction-channel graph can also be limited in the sense that one channel can be attached to at most one instruction. We show, for instance, that some VMs with a tree instruction or host graph and with some lower bounds on the number of hosts, instructions, and viruses can only compute finite sets. Lastly, we highlight a new (and better) characterisation of semilinear sets of numbers, in short $SLIN$, with virus machines. Our results on normal forms are then used to ask new questions regarding other normal forms and restrictions on VMs.
@@ 2.1 Virus Machines
Let a virus machine $\Pi$ of degree $(p,q)$ with $p,q\geq 0$ defined as:
$\Pi=(\Gamma,H,I,D_{H},D_{H},G_{C},n_{1},\dots,n_{p},i_{1},h_{out})$
@@
  • $H=\{h_{1},\dots,h_{p}\}$ is the ordered set of hosts, $h_{out}$ can be either in $H$ or not (for this work, we will suppose always $h_{out}\notin H$, $I=\{i_{1},\dots,i_{q}\}$ the ordered set of instructions.
  • $D_{H}=(H\cup\{h_{out}\},E_{H},w_{H})$ is the weighted and directed (WD) host graph, where the edges are called channels and $w_{H}:H\times H\cup\{h_{out}\}\rightarrow\mathbb{N}$.
  • $D_{I}=(I,E_{I},w_{I})$ is the WD instruction graph and $w_{I}:I\times I\rightarrow\{1,2\}$.
  • $G_{C}=(E_{H}\cup I,E_{I})$ is a unweighted bipartite graph called channel-instruction graph, where the partition associated is $\{E_{H}\cup I\}$.
@@
From a configuration $\mathcal{C}_{t}$, $\mathcal{C}_{t+1}$ is obtained as follows. The instruction that will be activated is $u_{t}$ if $u_{t}\in I$, otherwise $\mathcal{C}_{t}$ is a halting configuration. Let us suppose that $u_{t}\in I$ and that it is attached to the channel $(h_{j},h_{j^{\prime}})\in E_{H}$ with weight $w\in\mathbb{N}$, then the channel is opened and two possibilities holds:
  • If $a_{j,t}>0$, then there is virus transmission, that is, one virus is consumed from $h_{j}$ and is sent to the host $h_{j^{\prime}}$ replicated by $w$. The next activated instruction will follow the highest weight path in the instruction graph. In case the highest path is not unique, it is chosen nondeterministically. In case there is no possible path, then $u_{t+1}=\#$
  • If $a_{j,t}=0$, then there is no virus transmission and the next instruction follows the least weight path. For the other cases, it is analogous to the previous assumption.
This paper is focused on the computational power of VMs in the generating mode, and for that we fix the same notation as in the foundational paper [14]. Let
$NVM(p,q,n)$ be the family of sets of natural numbers generated by virus machines with at most $p$ hosts, $q$ instructions, and $n$ viruses in each host at any instant of the computation. For unbounded restrictions, they are replaced by *.
@@ 2.2 Topology
We say that $v_{1}$ is connected to $v_{n}$ if there is a path $w=(e_{1},\dots,e_{n-1})$, whose sequence of vertices is $(v_{1},\dots,v_{n})$. We denote by $V(v_{i})\subseteq V$ the subset of vertices that are connected by a path from $v_{i}$.
@@ 2.3 State-of-the-art
Figure 1: A virus machine generating $NFIN$ for $NVM(1,*,*)$.
Figure 2: A virus machine generating $NFIN$ for $NVM(*,*,1)$.
The state-of-the-art is presented in Table 1. The virus machines in the generating mode are Turing Universal; that is, they can generate recursively enumerable sets of numbers ($NRE$[1] for unbounded restrictions. This power is severely reduced when the last ingredient is reduced; more precisely, a semilinear set characterisation ($SLIN$) is proved for $NVM(*,*,2)$ [14]. From now on, not characterisations but inclusions have been proven, for finite sets ($NFIN$) they are contained in $NVM(1,*,*)$ and $NVM(*,*,1)$ [4]. Finally, the set of power of two numbers is contained in $NVM(2,7,*)$ [4].
@@
An interesting and natural question is can we further restrict or provide better lower bounds, for known results about VMs? That is, provide “better” characterisations of finite sets or even other families of sets such as the singleton sets, see, for instance, Table 1. As we focus on finite sets later, let us see the VMs used in [4] to generate finite sets. For $NVM(1,*,*)$ the VM presented in Figure 1, and for $NVM(*,*,1)$ the Figure 2. The corresponding lemmas were called (viruses) and (hosts), respectively, and we follow the same notation in this work.
@@ 3.1 Finite sets
$NFIN\subsetneq NVM(2,*,2)$.
@@ 3.2 Singleton Sets
    Let us start with the inclusion of the left side, let $\Gamma=\{v\}$ be a singleton set of natural number $v\in\mathbb{N}$, then it can be generated by the VM $\Pi_{sing_{1}}$ of degree $(1,1)$ depicted in Figure 5, the initial configuration is $\mathcal{C}_{0}=(1,i_{1},0)$ and in the following configuration, a virus is consumed and replicated by the weight of the arc, that is, $v$, and sent to the environment, leading to the halting configuration $\mathcal{C}_{1}=(0,\#,v)$. Thus, after one transition step, the set generated is $\{v\}$.
@@
    For reverse inclusion, suppose any VM with only one host and one virus: the host can only be attached to the environment, and let us fix that the weight of that channel is $w\in\mathbb{N}$. Thus, the only number generated is $w$ or none, depending on the instruction graph (if the computation halts or not). Thus, we generate a singleton set.
@@
    For the inclusion on the left side we can use the VM $\Pi_{sing_{1}}$ depicted in Figure 5 as it only has one instruction and the inclusion has already been proven.
    Let us focus on the inclusion of the right side. With only one instruction, there are two possibilities in the instruction graph:
@@ 3.3 Finite linear progressions
Let $NLinFIN=\bigcup_{x\in\mathbb{N}}\bigcup_{n\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}(\{x+n\cdot i:0\leq i\leq N\})\cup\{\ \emptyset\}$, the family of finite linear progressions. The following result holds.
@@
$NVM(p,2,*)=NLinFIN,$ for each $p\geq 2$.
@@
For the right-side inclusion, for any $x,n,N\in\mathbb{N}$, let us see that there exists a VM $\Pi_{Lin}$ of degree $(2,2)$ that generates the set $\{x+n\cdot i:0\leq i\leq N\}$. The virus machine can be depicted in Figure 6.
@@
For the inclusion on the left side, some previous considerations should be taken into account. First, let us fix that $I=\{i_{1},i_{2}\}$ is the set of instructions and $i_{1}$ is the initial instruction. We will consider the instruction graphs where there is at least one halting computation. On the other hand, we will consider those in which there is more than one computation, that is, there is a least one non-deterministic decision, otherwise only the singleton sets can be generated (which are included in $NLinFIN$).
@@ 3.4 Finite linear combinations
For each $w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}$. Finally, let $A=\{a_{w_{1},w_{2},N_{1},N_{2}}\}_{w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}},\\ B=\{b_{w_{1},w_{2},N_{1},N_{2}}\}_{w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}},$ be two family of sets of natural numbers, we define finite linear combinations $NCombFIN=\{\emptyset\}\cup A\cup B$.
@@
$NCombFIN=NVM(p,3,*),$ for each $p\geq 3$.
@@ 3.5 Discussion of old ingredients
In this subsection a brief discussion of the results obtained with the previous results is presented in Table 2.
@@
$NLinFIN$ (Proposition 3.7) $=$ 2 2 *
$NCombFIN$ (Proposition 3.9) $=$ 3 3 *
@@ 4 Novel results with new ingredients
For virus machines in generating mode, the notation used in previous works [4] was $NVM(p,q,r)$ where it denotes the family of natural number sets generated by virus machines with at most $p$ hosts, $q$ instructions and $r$ viruses at most at each instant in the computation. The computational completeness has been proven when these ingredients are unbounded [1], that is, $NRE=NVM(*,*,*)$, but the power decreases when one of them is bounded, for example, a characterisation of semilinear sets has been proven in [14], that is, $SLIN=NVM(*,*,r)$ for each $r\geq 2$. However, the three ingredients mentioned seem to be poor information. We propose the following notation:
$NVM_{\beta}(h_{p},i_{q},nvh_{r},wc_{s},outd_{t},\alpha_{\hbar}^{u},\alpha_{\iota}^{v}),$ (1)
@@
First, we believe that the instruction-channel graphs have something to say in the sense of computational power, that is why we fix bijection that will show a new frontier of computational power when we combine it with other ingredients.
Secondly, fixing the directed graphs as trees is a first intention to approach the topology structure of the graphs, fixing them as trees (when $u$ or $v$ are zero) will allow us to fix characterisations to finite sets, as we will see later. Nevertheless, we believe that not only whether or not there is a loop is important, but also the size of the loop. The relation between these sizes and the power of the devices will have interesting results. Note that $u=1$ is the same as $u=0$ as there are no self-arcs in the host graph.
In addition, the weight of the arcs is our unique way to increase the amount of viruses, fixing this weight to one where the power falls substantially, but if it is $2$ we can get novel universality results.
Lastly, the out-degree of the hosts, we believe that it is also crucial in the restrictions of the computing power, it will be shown in combination with other ingredients.

### v2

@@ Abstract
In the present work, we further study the computational power of virus machines (VMs in short). VMs provide a computing paradigm inspired by the transmission and replication networks of viruses. VMs consist of process units (called hosts) structured by a directed graph whose arcs are called channels and an instruction graph that controls the transmissions of virus objects among hosts. The present work complements our understanding of the computing power of VMs by introducing normal forms; these expressions restrict the features in a given computing model. Some of the features that we restrict in our normal forms include (a) the number of hosts, (b) the number of instructions, and (c) the number of virus objects in each host. After we recall some known results on the computing power of VMs we give our series of normal forms, such as the size of the loops in the network, proving new characterisations of family of sets, such as finite sets, semilinear sets, or recursively enumerable sets ($NRE$).
@@ 1 Introduction
This paper contributes the following to the study of virus machines and their computing power. Normal forms, some of which are optimal bounds, for VMs are provided in the following sense: (a) providing characterisations (previously were inclusions) for generating families of finite sets; (b) showing new characterisations for finite sets of numbers using restrictions on the number of required hosts, instructions, or viruses; and (c) new characterisations are also given for singleton sets of numbers and some linear progressions or combinations. The previous work was focused on the restriction of only thre ingredients: hosts, instructions, and number of virueses. In this work, new restrictions are considered: limiting or not the host or instruction graphs to be a tree graph, that is, an acylcic directed graph. Moreover, the instruction-channel graph can also be limited in the sense that one channel can be attached to at most one instruction. We show, for instance, that some VMs with a tree instruction or host graph and with some lower bounds on the number of hosts, instructions, and viruses can only compute finite sets. Lastly, we highlight a new (and better) characterisation of semilinear sets of numbers, in short $SLIN$, with virus machines. Our results on normal forms are then used to ask new questions regarding other normal forms and restrictions on VMs.
@@ 2.1 Virus Machines
Let a virus machine $\Pi$ of degree $(p,q)$ with $p,q\geq 0$ the number of hosts and instructions respectively, defined as:
$\Pi=(\Gamma,H,I,D_{H},D_{I},G_{C},n_{1},\dots,n_{p},i_{1},h_{out})$
@@
  • $H=\{h_{1},\dots,h_{p}\}$ is the ordered set of hosts, $h_{out}$ can be either in $H$ or not (for this work, we will suppose always $h_{out}=h_{0}\notin H$, $I=\{i_{1},\dots,i_{q}\}$ the ordered set of instructions.
  • $D_{H}=(H\cup\{h_{out}\},E_{H},w_{H})$ is the weighted and directed (WD) host graph, where the edges are called channels and $w_{H}:E_{H}\cup\{h_{out}\}\rightarrow\mathbb{N}$.
  • $D_{I}=(I,E_{I},w_{I})$ is the WD instruction graph and $w_{I}:E_{I}\rightarrow\{1,2\}$.
  • $G_{C}=(E_{H}\cup I,E_{I})$ is a unweighted bipartite graph called channel-instruction graph, where the disjoint partition associated is $\{E_{H}\cup I\}$, although a channel can be attached to more than one instruction, each instruction is attached to only one channel at most.
@@
From a configuration $\mathcal{C}_{t}$, $\mathcal{C}_{t+1}$ is obtained as follows. The instruction that will be activated is $u_{t}$ if $u_{t}\in I$, otherwise $\mathcal{C}_{t}$ is a halting configuration. Let us suppose that $u_{t}\in I$ and that it is attached to the channel $(h_{j},h_{j^{\prime}})\in E_{H}$ with weight $w_{H}((h_{j},h_{j^{\prime}}))=w\in\mathbb{N}$, then the channel is opened and two possibilities holds:
  • If $a_{j,t}>0$, then there is virus transmission, that is, one virus is consumed from $h_{j}$ and is sent to the host $h_{j^{\prime}}$ replicated by $w$. The next activated instruction will follow the highest weight path from $u_{t}$ to just the next instruction on the instruction graph. In case the highest path is not unique, it is chosen nondeterministically. In case there is no possible path, then $u_{t+1}=\#$
  • If $a_{j,t}=0$, then there is no virus transmission and the next instruction follows the least weight path from $u_{t}$ to just the next instruction. For the other cases, it is analogous to the previous assumption.
This paper is focused on the computational power of VMs in the generating mode, that is the numbers that can be generated in all the possible computations, and for that we fix the same notation as in the foundational paper [14] which will be called old ingredients. Let $\mathrm{NVM}(p,q,n)$ be the family of sets of natural numbers generated by virus machines with at most $p$ hosts, $q$ instructions, and $n$ viruses in each host at any instant of the computation. For unbounded restrictions, they are replaced by *.
2.1.1 Explicit Example
To clarify the behavior of these devices, the generating mode, and their visual representation, let us see an explicit example. Let $\Pi_{ex}$ be the VM of degree $(2,4)$ in the generating mode and visually described in Figure 1. The squares represent the host units with $h_{0}$ being the environment, and the double arrows the arcs of the host graph, for simplicity, if the weight $1$ then the number is omitted in the representation. The blue dots are the instructions units and the vertices in the instruction graph. Lastly, the red dashed lines are the edges of the instruction-channel graph.
Figure 1: The virus machine $\Pi_{ex}$.
The computation works as follows, the initial configuration is $\mathcal{C}_{0}=(2,2,i_{1},0)$. The instruction $i_{1}$ is connected to the channel $(h_{1},h_{0}),$ as the host $h_{1}$ is not empty, there is virus transmission and the highest weight path is followed, that is going again to $i_{1}$ with configuration $\mathcal{C}_{1}=(1,2,i_{1},1)$. The process is analogous and configuration $\mathcal{C}_{2}=(0,2,i_{1},2)$ is reached. From here, as host $h_{1}$ is empty, there is no virus transmission and the least weight path is followed, having the configuration $\mathcal{C}_{3}=(0,2,i_{2},2)$. Instruction $i_{2}$ opens the channel $(h_{2},h_{1})$ so there is virus transmission, but now the highest weight path is not unique, having two possible configurations depending on the instruction that is chosen:
  • If instruction $i_{3}$ is chosen, then the configuration is $\mathcal{C}_{4}=(1,1,i_{3},2)$. That instruction is connected to the channel $(h_{2},h_{0})$ with weight $2$, thus the virus sent from host $h_{2}$ is replicated by $2$ and sent to the output region. As there is no possible path from the instruction graph, a halting configuration is reached $\mathcal{C}_{5}=(1,1,\#,4)$. This means that the number $4$ has been generated.
  • If instruction $i_{4}$ is chosen, then the configuration is $\mathcal{C}_{4}=(1,1,i_{4},2)$. As that instruction is not attached to any channel and there is no possible path in the instruction graph, a halting configuration is reached $\mathcal{C}_{5}=(1,1,\#,2)$. This means that the number $2$ has been generated.
With all the possible computations checked, we can state that the VM $\Pi_{ex}$ generates the set $\{2,4\}$.
@@ 2.2 Topology
We say that $v_{1}$ is connected or attached to $v_{n}$ if there is a path $w=(e_{1},\dots,e_{n-1})$, whose sequence of vertices is $(v_{1},\dots,v_{n})$. We denote by $V(v_{i})\subseteq V$ the subset of vertices that are connected by a path from $v_{i}$.
@@ 2.3 State-of-the-art
Figure 2: A virus machine generating $\mathrm{NFIN}$ for $\mathrm{NVM}(1,*,*)$.
Figure 3: A virus machine generating $\mathrm{NFIN}$ for $\mathrm{NVM}(*,*,1)$.
The state-of-the-art is presented in Table 1. The virus machines in the generating mode are Turing Universal; that is, they can generate recursively enumerable sets of numbers ($NRE$[1] for unbounded restrictions. This power is severely reduced when the last ingredient is reduced; more precisely, a semilinear set characterisation ($SLIN$) is proved for $\mathrm{NVM}(*,*,2)$ [14]. From now on, not characterisations but inclusions have been proven, for finite sets ($\mathrm{NFIN}$) they are contained in $\mathrm{NVM}(1,*,*)$ and $\mathrm{NVM}(*,*,1)$ [4]. Finally, the set of power of two numbers is contained in $\mathrm{NVM}(2,7,*)$ [4].
@@
An interesting and natural question is can we further restrict or provide better lower bounds, for known results about VMs? That is, provide “better” characterisations of finite sets or even other families of sets such as the singleton sets, see, for instance, Table 1. As we focus on finite sets later, let us see the VMs used in [4] to generate finite sets. For $\mathrm{NVM}(1,*,*)$ the VM presented in Figure 2, and for $\mathrm{NVM}(*,*,1)$ the Figure 3. The corresponding lemmas were called (viruses) and (hosts), respectively, and we follow the same notation in this work.
@@ 3 Novel results with old ingredients
Having set the state-of-the-art of VM in generating mode, let us see new results by the boundary of those old ingredients: the hosts, the instructions, and the viruses during the computation. This section will start with finite sets, just to complete the previous results, and then it is organised from weak to stronger power (in the sense of bigger sets), in the end a table summarising all the novel results is presented.
@@ 3.1 Finite sets
$\mathrm{NFIN}\subsetneq\mathrm{NVM}(2,*,2)$.
@@ 3.2 Singleton Sets
    Let $\Gamma=\{v\}$ be a singleton set of natural number $v\in\mathbb{N}$, then it can be generated by the VM $\Pi_{sing_{1}}$ of degree $(1,1)$ depicted in Figure 6, the initial configuration is $\mathcal{C}_{0}=(1,i_{1},0)$ and in the following configuration, a virus is consumed and replicated by the weight of the arc, that is, $v$, and sent to the environment, leading to the halting configuration $\mathcal{C}_{1}=(0,\#,v)$. Thus, after one transition step, the set generated is $\{v\}$.
@@
    superset-of-or-equals Suppose any VM with only one host and one virus: the host can only be attached to the environment, and let us fix that the weight of that channel is $w\in\mathbb{N}$. Thus, the only number generated is $w$ or none, depending on the instruction graph (if the computation halts or not). Thus, we generate a singleton set.
@@
    Let us use the VM $\Pi_{sing_{1}}$ depicted in Figure 6 as it only has one instruction and the inclusion has already been proven.
    superset-of-or-equals With only one instruction, there are two possibilities in the instruction graph:
@@ 3.3 Finite linear progressions
Let $\mathrm{NLinFIN}=\bigcup_{x\in\mathbb{N}}\bigcup_{n\in\mathbb{N}}\bigcup_{N\in\mathbb{N}}(\{x+n\cdot i:0\leq i\leq N\})\cup\{\ \emptyset\}$, the family of finite linear progressions. The following result holds.
@@
$\mathrm{NVM}(p,2,*)=\mathrm{NLinFIN},$ for each $p\geq 2$.
@@
superset-of-or-equals For any $x,n,N\in\mathbb{N}$, let us see that there exists a VM $\Pi_{Lin}$ of degree $(2,2)$ that generates the set $\{x+n\cdot i:0\leq i\leq N\}$. The virus machine can be depicted in Figure 7.
@@
Some previous considerations should be taken into account. First, let us fix that $I=\{i_{1},i_{2}\}$ is the set of instructions and $i_{1}$ is the initial instruction. We will consider the instruction graphs where there is at least one halting computation. On the other hand, we will consider those in which there is more than one computation, that is, there is a least one non-deterministic decision, otherwise only the singleton sets can be generated (which are included in $\mathrm{NLinFIN}$).
@@ 3.4 Finite linear combinations
For each $w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}$. Finally, letting $A=\{a_{w_{1},w_{2},N_{1},N_{2}}\}_{w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}},\\ B=\{b_{w_{1},w_{2},N_{1},N_{2}}\}_{w_{1},w_{2},N_{1},N_{2}\in\mathbb{N}},$ be two families of sets of natural numbers, we define finite linear combinations $\mathrm{NCombFIN}=\{\emptyset\}\cup A\cup B$.
@@
$\mathrm{NCombFIN}=\mathrm{NVM}(p,3,*),$ for each $p\geq 3$.
@@ 3.5 Discussion of old ingredients
In this subsection a brief discussion of the results obtained with the previous results is presented in Table 2. Here some new characterisations have been proven below the finite sets, such as the singleton sets, the $\mathrm{NLinFIN}$ and the $\mathrm{NCombFIN}$. This was possible due to the limit of the number of instructions. Lastly, a new inclusion of the finite sets (and even more, strict inclusion) has been stated.
@@
$\mathrm{NLinFIN}$ (Proposition 3.7) $=$ 2 2 *
$\mathrm{NCombFIN}$ (Proposition 3.9) $=$ 3 3 *
@@ 4 Novel results with new ingredients
For virus machines in generating mode, the notation used in previous works [4] was $\mathrm{NVM}(p,q,r)$ where it denotes the family of natural number sets generated by virus machines with at most $p$ hosts, $q$ instructions and $r$ viruses at each instant in the computation. The computational completeness has been proven when these ingredients are unbounded [1], that is, $\mathrm{NRE}=\mathrm{NVM}(*,*,*)$, but the power decreases when one of them is bounded, for example, a characterisation of semilinear sets has been proven in [14], that is, $\mathrm{SLIN}=\mathrm{NVM}(*,*,r)$ for each $r\geq 2$. However, the three ingredients mentioned seem to be poor information. We propose the following notation:
$\mathrm{NVM}_{\beta}(h_{p},i_{q},nvh_{r},wc_{s},outd_{t},\alpha_{\hbar}^{u},\alpha_{\iota}^{v}),$ (1)
@@
First, we believe that the instruction-channel graphs have something to say in the sense of computational power, that is why we fix bijection that will show a new frontier of computational power when we combine it with other ingredients for instance in theorems 4.44.9 and  4.11.
Secondly, fixing the directed graphs as trees is a first intention to approach the topology structure of the graphs, fixing them as trees (when $u$ or $v$ are zero) will allow us to fix characterisations to finite sets (such as Corollary 4.3 or Theorem 4.4). Nevertheless, we believe that not only whether or not there is a loop is important, but also the size of the loop. The relation between these sizes and the power of the devices will have interesting results, for instance theorems  4.94.114.12 and 4.18, and Lemma 3.1. Note that $u=1$ is the same as $u=0$ as there are no self-arcs in the host graph.
In addition, the weight of the arcs is our unique way to increase the amount of viruses, fixing this weight to one where the power falls substantially, but if it is $2$ we can get novel universality results (see Theorem 4.18).
Lastly, the out-degree of the hosts, we believe that it is also crucial in the restrictions of the computing power, it will be shown in combination with other ingredients (see all the mentioned theorems and lemmas).